OpenStudy (anonymous):

Find the sum:

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} 5 (1/4)^n\]

OpenStudy (anonymous):

What type of math is this lol

OpenStudy (anonymous):

calc II

OpenStudy (anonymous):

great. I'm doing precalc :P

OpenStudy (anonymous):

calc II is basically integrals and sequences, no fun.

OpenStudy (anonymous):

simple a=5(1/4)^1=5/4 r=5(1/4)^2/5(1/4)^1=5*4/16*5 =1/4 now sum to infinity=a/1-r=(5/4)/(1-1/4) =5*4/3*4=5/3